OpenStudy (turingtest):

I was trying to explain to someone why a+b(c) is not equal to (a+b)(c). Does anybody have an idea how to convince somebody of this through common sense?

OpenStudy (anonymous):

a+b(c)=a+bc

OpenStudy (anonymous):

(a+b)(c)=ac+bc

OpenStudy (anonymous):

tell him to pick a book and learn the distributive law

OpenStudy (turingtest):

I was trying to teach my Spanish teacher this concept, I doubt she wants to pick up a book about algebraic properties and memorize them.

OpenStudy (jamesj):

Use a counterexample. If a = b = 1, c = 2, then a + b(c) = 1 + 2 = 3 but (a + b)c = 2x2 = 4

OpenStudy (anonymous):

if he's very stubborn plug in some values and work them out!!!

OpenStudy (turingtest):

Yeah James, I did that, but all she could reply to that was "so I have to multiply first?" I don't think it gave her an intuitive understanding of the concept.

OpenStudy (anonymous):

your spanish teacher? jesus christ... are these the people teaching nowadays?

OpenStudy (turingtest):

LOL I think she was doing her daughters homework. I don't want to say she's incapable of understanding this kind of thing, but I guess it seems that way.

OpenStudy (turingtest):

oh well, I guess some people just aren't cut out for this kind of thing...

OpenStudy (anonymous):

TuringTest i need your help with another ridiculous question lol

OpenStudy (turingtest):

sure, did you post it yet?

OpenStudy (anonymous):

Yeah i did

OpenStudy (turingtest):

send me the link, yeah?

OpenStudy (turingtest):

oh I found it...

OpenStudy (anonymous):

Ok thank you